Given function p and t defined by the table below
 x,  -2, -1,  0,  1,  2
p(x), 4,  1,  0,  1,  4
t(x), 3,  2,  1, -2, -3

Find
1. t(p(-1))
2. (t∘p)(0)
3  (p∘t)(-1)
4. t(t(0))
5. P(t(2))

Answer by Edwin McCravy(20056)

Given function p and f defined by the table below:
  x  | -2, -1,  0,  1,  2
-------------------------
p(x) |  4,  1,  0,  1,  4
t(x) |  3,  2,  1, -2, -3


1. t(p(-1))           = t(1) = -2

2. (t∘p)(0) = t(p(0)) = t(0) =  1 

3  (p∘t)(-1) =  ?     =  ?   =  ?

4. t(t(0))            =  ?   =  ?

5. P(t(2))            =  ?   =  ?

You do 3,4, and 5.  It's just a matter of looking in the table,
and doing exactly what I did in the first two.

Edwin